Evaluate. 4 · 2^2 - 2 · 8
Answer by fractalier(2101)

4 · 2^2 - 2 · 8
We use the order of operations, shown by the memory device PEMDAS; so exponents go first...then multiplying, then subtracting...
4 · 2^2 - 2 · 8 =
4 • 4 - 2 • 8 =
16 - 16 =
0